please help me solve the following problem.
7/8x - 1/4 + 3/4x = 1/16 + x

Answer by stanbon(75887)

7/8x - 1/4 + 3/4x = 1/16 + x
-----
Multiply thru by 16
14x - 4 + 12x = 1 + 16x
26x-4 = 1 + 16x
10x = 5
x = 1/2
==============
Cheers,
Stan H.
==============